RunPlan20161128 » History » Revision 62
« Previous |
Revision 62/87
(diff)
| Next »
MOM, Mice Operations Manager, 10 December 2016 14:17
RunPlan20161128 to Dec 16th 2016¶
Scattering data with field, backup plan for straight tracks.
Daily schedule from 28/Nov to 16/Dec.¶
We expect to be running with 2V beamloss. Note Decay Solenoid is ON. ISIS is at 40+10 and 700 MeV
Note:
Date | Dec 4 | Dec. 5 | Dec. 6 | Dec. 7 | Dec. 8 | Dec. 9 | Dec 10 |
MOM | M. Uchida | M. Uchida | M. Uchida | M. Uchida | M. Uchida | M. Uchida | M. Uchida |
BLOC | R.Gamet | R.Gamet | R.Gamet | R.Gamet | J.Greis | J.Greis | J.Greis |
SOC | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram |
Lead shifter (08:00-16:00) | R.Gamet | R.Gamet | R.Gamet | R.Gamet | S.Wilbur | S.Wilbur | S.Wilbur |
Shifter 2 (08:00-16:00) | D.Bowring | D.Bowring | D.Bowring | D.Bowring | J.Greis | J.Greis | J.Greis |
Lead shifter (16:00-00:00) | P.Soler | P.Soler | P.Soler | P.Soler | D.Neuffer | D.Neuffer | D.Neuffer |
Shifter 2 (16:00-00:00) | D.Maletic | D.Maletic | D.Maletic | D.Maletic | D.Maletic | D.Maletic | D.Maletic |
Lead Shifter (00:00-08:00) | M.Courthold | M.Courthold | P.Kyberd | P.Kyberd | A.Bross | A.Bross | A.Bross |
Shifter (00:00-08:00) | B.Freemire | B.Freemire | P.Hanlet | P.Hanlet | B.Freemire | B.Freemire | B.Freemire |
Additional Personnel | |||||||
Notes | |||||||
Date | Dec 11 | Dec. 12 | Dec. 13 | Dec. 14 | Dec. 15 | Dec. 16 |
MOM | M. Uchida | M. Uchida | M. Uchida | M. Uchida | M. Uchida | M. Uchida |
BLOC | J.Greis | J.Greis | Y.Song | Y.Song | Y.Song | Y.Song |
SOC | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram |
Lead shifter (08:00-16:00) | S.Wilbur | S.Wilbur | S.Boyd | S.Boyd | S.Boyd | |
Shifter 2 (08:00-16:00) | J.Greis | J.Greis | B.Freemire | B.Freemire | B.Freemire | |
Lead shifter (16:00-00:00) | D.Neuffer | D.Neuffer | M.Bonesini | M.Bonesini | M.Bonesini | |
Shifter 2 (16:00-00:00) | B.Freemire | B.Freemire | L.Tortora | L.Tortora | L.Tortora | |
Lead Shifter (00:00-08:00) | A.Bross | A.Bross | D.Neuffer | D.Neuffer | D.Neuffer | |
Shifter (00:00-08:00) | D.Maletic | D.Maletic | Y.Song | Y.Song | Y.Song | |
Additional Personnel | ||||||
Notes | ||||||
Plan from 9 Dec:¶
Take H10a-c data in two hour chunks until complete.
Then Take H9d data in two hour chunks until complete
Data to take¶
Field On Data¶
Task H9 - Emittance Reduction!!¶
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs | Final Event Tally (k) |
H9d | 200 | 2e6 TOF1 | Diff Mask 8 200+M3-Test1 | 2016-04-1.7 |
| RUN Settings | Diffuser | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 |
| | | mm | A | A | A | A | A | A | A | A | A | A | A | A |
| H9d | Mask 8 (Iris 4 closed 1-3 open) | 29 | | | | | ----| | | | | | | |
h4. Task H10 - Magnet alignment
No beamline tags exist. All H10 runs will be made with the magnets manually set to match the beamline tag but with the iris 4 closed and in the comment write eg. "Diff Mask 8 -140+M3_Test2 settings"
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs |
H10a | 140 | 7e5 TOF1 | Diff Mask 8 as 3-140+M3_Test2 | 2016-04-1.7 | |
H10b | 170 | 7e5 TOF1 | Diff Mask 8 as 3-170+M3-Test1 | 2016-04-1.7 | |
H10c | 240 | 7e5 TOF1 | Diff Mask 8 as 3-240+M3_Test1 | 2016-04-1.7 |
Straight track data (fallback)¶
Task S2: LiH scattering data.¶
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs |
S2a | 140 | 1M @ TOF2 | 3-140+M3-Test2 | ||
S2b | 200 | ? | 6-200+M0 |
Task S2a - 10M triggers at TOF1, 33 triggers/spill = 90 Hours data taking (15 shifts). Lots, but will make a
start anyway in case there is opportunity in later data taking for more running.
Finished data sets:¶
h4. Task H4 - Emittance Reduction!!
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs | Final event tally |
H4a | 140 | 1e6 TOF1 | 3-140+M3-Test2 | 2016-04-01.2 | 1749 | |
H4b | 140 | 1e6 TOF1 | 6-140+M3-Test2 | 2016-04-01.2 | 1694 | |
H4c | 140 | 2e6 TOF1 | 10-140+M3-Test3 | 2016-04-01.2 | 2909 |
h4. Task H5 - Magnet alignment (DONE)
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs | Final Event tally |
H5a | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 9.55e6 | |
H5b | 200 | 5e5 TOF1 | 3-200+M3_Test1 | 2016-04-01.2 | 8.92e6 | |
H5c | 240 | 5e5 TOF1 | 3-240+M3_Test1 | 2016-04-01.2 | 7.93e6 |
h4. Task H6 - Diffuser Scan
The beamline and cooling channel settings remain the same and the diffuser mus be changed manually from the diffuser GUI for each new setting! Be sure to do this AFTER entering the beamline optics, or it will reset the diffuser. (Please note there is no run H6d.
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Diffuser Setting | Closed Iris | Similar runs | Final Event Tally |
H6a | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 1 | 1 | 8498 | 320 |
H6b | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 2 | 2 | 8497 | 300 |
H6c | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 4 | 3 | 8495-6 | 300 |
H6e | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 6 | 2+3 | 8499 | 300 |
H6f | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 7 | 1+2+3 | 8502 | 301 |
H6g | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 8 | 4 | 8494 | 300 |
H6h | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 9 | 1+4 | 8500 | 300 |
H6i | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 12 | 3+4 | 8501 | 300 |
H6j | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-01.2 | 15 | 1+2+3+4 | 8503 | 300 |
h4. Task H7 - Emittance Reduction!!(DONE)
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs | final event tally |
H7a | 140 | 1e6 TOF1 | 3-140+M3-Test2 | 2016-04-1.5 | H4a | 1494k |
H7b | 140 | 1e6 TOF1 | 6-140+M3-Test2 | 2016-04-1.5 | H4b | 1432k |
H7c | 140 | 2e6 TOF1 | 10-140+M3-Test3 | 2016-04-1.5 | H4c | 2741k |
h4. Task H8 - Magnet alignment (DONE)
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs | final event tally |
H8a | 170 | 5e5 TOF1 | 3-170+M3-Test1 | 2016-04-1.5 | H5a | 668k |
H8b | 200 | 5e5 TOF1 | 3-200+M3_Test1 | 2016-04-1.5 | H5b | 667k |
H8c | 240 | 5e5 TOF1 | 3-240+M3_Test1 | 2016-04-1.5 | H5c | 665 |
h4. Task H9 - Emittance Reduction!! (DONE)
Task | Momentum | # Triggers | Beamline Tag | CoolingChannel Tag | Similar runs | Final Event Tally (k) |
H9a | 200 | 1e6 TOF1 | 3-200+M3_Test1 | 2016-04-1.7 | H4a H7a | 1208 |
H9b | 200 | 1e6 TOF1 | 6-200+M3-Test1 | 2016-04-1.7 | H4b H7b | 1000 |
H9c | 200 | 2e6 TOF1 | 10-200+M3-Test1 | 2016-04-1.7 | H4c H7c | 3817 |
Previous Shifts¶
Date | Nov 28 | Nov. 29 | Nov. 30 | Dec. 1 | Dec. 2 | Dec. 3 |
MOM | E. Overton | M. Uchida | M. Uchida | M. Uchida | M. Uchida | M. Uchida |
BLOC | Y.Song | Y.Song | Y.Song | Y.Song | Y.Song | R.Gamet |
SOC | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram |
shifter 2 (08:00-16:00) | Y.Song | Y.Song | Y.Song | Y.Song | Y.Song | D.Bowring |
Lead Shifter(08:00-16:00) | P. Franchini | P. Franchini | P. Franchini | P. Franchini | P. Franchini | P.Soler |
Lead shifter (16:00-00:00) | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | R.Gamet |
Shifter 2 (16:00-00:00) | B.Freemire | B.Freemire | B.Freemire | B.Freemire | B.Freemire | D.Maletic |
Lead Shifter (00:00-08:00) | G. Vankova | G. Vankova | G. Vankova | G. Vankova | G. Vankova | B.Freemire 4-8 |
Shifter (00:00-08:00) | J. Langlands | J. Langlands | J. Langlands | J. Langlands | J. Langlands | P.Hanlet 12-4 |
Additional Personnel | M.Courthold | |||||
Notes |
Updated by MOM, Mice Operations Manager almost 7 years ago · 62 revisions